# Combining Constrained and Unconstrained Decoding via Boosting: BoostCD and Its Application to Information Extraction

### v1

@@ Abstract
Marija Šakota1  Robert West1,2 1EPFL  2Microsoft {marija.sakota, robert.west}@epfl.ch
@@ 2 Method: BoostCD
Language models trained for structured NLP tasks in a supervised manner can often perform reasonably well even without the constraints imposed, but the constraints are still required to guarantee 100% valid generations, and they can steer the model to pick the one correct output when multiple outputs might seem plausible a priori (e.g., when an entity has multiple aliases). However, when the constraints require altering the unconstrained output significantly (e.g., when an entity generated in unconstrained mode is not present in the KB), performance can suffer from imposing the constraints.
@@
(i) Phase 1: We use a base model $M$, trained in an autoregressive manner on $(x,y)$ pairs, to make two parallel passes. In one, we let the model generate in an unconstrained manner: by sending input text $x$ to the model $M$ without imposing any constraints, we obtain the output $y_{u}$. In the second, we generate by imposing constraints: by providing the input text $x$ and using $M$ with constrained decoding, we obtain the output $y_{c}$.
@@
During the inference, we repeat the steps from both phases: (1) we make two parallel passes with the base model $M$ to generate constrained and unconstrained predictions ($\hat{y_{c}}$ and $\hat{y_{u}}$) and (2) we send $(x,\hat{y_{u}},\hat{y_{c}})$ to the input of the boosted model $M_{b}$ to make a final prediction $\hat{y}$. This prediction can be made with our without constrained decoding.
In the following sections, we apply BoostCD to the cIE task by curating the data and modeling to fit its needs. We emphasize that this paradigm can be used for other structured tasks, with adaptations of the data and modeling. Note that we use only one step of boosting in our pipeline, although in principle there is nothing that restricts this pipeline to one iteration only. For our setting, we found one step to be sufficient, but for other applications, it is possible to explore multiple iterations of the same algorithm.
@@ 3.1 Data
To train the base model, we need data that is exhaustive, i.e. the input is fully expressible under constraints. For cIE, the base model should extract all the facts present in the text, regardless of constraints (i.e. perform open information extraction).For the boosted model, we can simulate the setting in which some samples express entities in the text that do not exist in the KB. For a fraction of the data we randomly remove some entities from the KB making it impossible for a base model to generate them in the constrained setting. We also remove these entities from the target triplet set by removing each triplet that contains the entity in question. By providing these samples during training, we let the model learn what happens when the entity in text is not present in the KB and hopefully bring it closer to generating the correct output.
@@ 3.3 DPO finetuning
As we currently do not have access to a well-aligned dataset for cIE that is made on real-world data, the process of training base and boosted models is done with synthetic data that might not highly resemble natural text. As a consequence, this might hinder the performance of our model in the wild. In an attempt to overcome this, we propose to tune the model with Direct Preference Optimization (DPO) Rafailov et al. (2023), using data more similar to the real-world one.
@@ 4 Evaluation setup
Data. For training the base model $M$, we use a subset of 300K samples from the train split of Wiki-cIE Code used for training SynthIE models (see Appendix B for details). The boosted model $M_{b}$ was trained on an additional 100K samples from the same dataset. We also train a SynthIE model on all 400k samples for a fair comparison. 100K samples used for training the boosted model have been altered as explained in Sec. 3.1, and 40% of the randomly chosen samples have been altered. For each altered sample, up to three entities were removed, uniformly. The validation and test data were crafted in the same way, each being a subset of the corresponding Wiki-cIE Code of 10K samples. Wiki-cIE Code is imperfect as it does not have the same properties as the real-world text, but can demonstrate the abilities of this training technique effectively.
@@ Appendix B Wiki-cIE Code
[…]
@@ Appendix E Implementation
As mentioned in Sec. 3.2, BoostIE uses two FlanT5 models. For both models, we use ’google/flan-t5-base’ version [^3] https://huggingface.co/google/flan-t5-base, which has $\sim$250M parameters. The models were trained using the Adam optimizer with a learning rate of 3e-4, 0.1 gradient clipping on the Euclidean norm, and a weight decay of 0.05. They were trained with batch size 80, for a maximum of 10K steps. We used a polynomial learning rate scheduler with 1000 warm-up steps and a final learning rate of 3e-05. All the experiments were run on a single NVIDIA Titan X Maxwell 12GB GPU, taking around 24h for the training of the base model, and around 16h for boosted model. The DPO finetuning was done on the same machine, using learning rate 5e-5, batch size 2, $\beta$ 0.1 and running it for 5 epochs, taking around 20min to finetune. During inference, we run all our models with 10 beams.

### v2

@@ Abstract
Marija Šakota   Robert West EPFL, Lausanne, Switzerland {marija.sakota, robert.west}@epfl.ch
@@ 2 Method: BoostCD
Language models trained for structured NLP tasks in a supervised manner can often perform reasonably well even without the constraints imposed, but the constraints are still required to guarantee 100% valid generations, and they can steer the model to pick the one correct output when multiple outputs might seem plausible a priori (e.g., when an entity has multiple aliases). However, when the constraints require altering the unconstrained output significantly (e.g., when an entity generated in unconstrained mode is not present in the KB), performance can suffer from imposing the constraints (for a more formal evaluation of this, see analysis in Appendix H.1).
@@
(i) Phase 1: We use a base model $M$, trained in an autoregressive manner on $(x,y)$ pairs, to make two parallel passes. In one pass, we let the model generate in an unconstrained manner: by sending input text $x$ to the model $M$ without imposing any constraints, we obtain the output $y_{u}$. In the other pass, we generate by imposing constraints: by providing the input text $x$ and using $M$ with constrained decoding, we obtain the output $y_{c}$.
@@
[…]
In the following sections, we apply BoostCD to the cIE task by curating the data and modeling to fit its needs. We emphasize that this paradigm can be used for other structured tasks, with adaptations of the data and modeling. Also note that we use only one step of boosting in our pipeline, although in principle there is nothing that restricts this pipeline to one iteration only. For our setting, we found one step to be sufficient, but for other applications, it is possible to explore multiple iterations of the same algorithm.
@@ 3.1 Data
To train the base model, we need data that is exhaustive, i.e. the input is fully expressible under constraints. In other words, there should be no facts in the text that we cannot express with entities and relations from the KB. We need this to train a model that is exhaustive. For cIE, this means that the base model should extract all the facts present in the text, regardless of constraints (i.e., perform open information extraction). If this was not the case, the model will likely learn to drop random triplets, and hence the performance would drop. For the boosted model, we can simulate the setting in which some samples express entities in the text that do not exist in the KB. For a fraction of the data we randomly remove some entities from the KB making it impossible for a base model to generate them in the constrained setting. We also remove these entities from the target triplet set by removing each triplet that contains the entity in question. By providing these samples during training, we let the model learn what happens when the entity in text is not present in the KB and hopefully bring it closer to generating the correct output.
@@ 3.3 DPO finetuning
As we currently do not have access to a well-aligned dataset for cIE that is made on real-world data (see Sec. 4), the process of training base and boosted models is done with synthetic data that might not highly resemble natural text. As a consequence, this might hinder the performance of our model in the wild. In an attempt to overcome this, we propose to tune the model with Direct Preference Optimization (DPO) Rafailov et al. (2023), using data more similar to the real-world one.
@@ 4 Evaluation setup
Data. For training the base model $M$, we use a subset of 300K samples from the train split of Wiki-cIE Code used for training SynthIE models (see Appendix B for details). This is a synthetic dataset and we use it because there is no real-text dataset with inputs that are fully expressible under KB constraints (see Sec. 3.1 for a further explanation of this requirement). The boosted model $M_{b}$ was trained on an additional 100K samples from the same dataset. We also train a SynthIE model on all 400k samples for a fair comparison. 100K samples used for training the boosted model have been altered as explained in Sec. 3.1, and 40% of the randomly chosen samples have been altered. For each altered sample, up to three entities were removed, uniformly. The validation and test data were crafted in the same way, each being a subset of the corresponding Wiki-cIE Code of 10K samples. Wiki-cIE Code is imperfect as it does not have the same properties as the real-world text, but can demonstrate the abilities of this training technique effectively.
@@ Appendix B Wiki-cIE Code
Wiki-cIE Code is a fully synthetic dataset introduced by Josifoski et al. (2023). It was used for training the range of SynthIE models. The dataset consists of around 1.8M training data samples, 10K validation, and 50K test samples generated by the now discontinued OpenAI model, code-davinci-002. The data was synthetically made, starting from sampling triplet sets. Triplet sets are generated by a biased random walk on a subset of the Wikidata knowledge graph Vrandečić (2012). Text that corresponds to these triplets was then generated by an LLM. Each text sample was generated by providing a triplet set and asking the LLM to write the text that only expresses those triplets. This way, an exhaustive, high-quality data was made. The main disadvantage of this dataset is the fact that the text does not resemble real text, as it is very clean and does not contain a lot of details.
@@ Appendix E Implementation
As mentioned in Sec. 3.2, BoostIE uses two FlanT5 models. For both models, we use ’google/flan-t5-base’ version [^3] https://huggingface.co/google/flan-t5-base, which has $\sim$250M parameters. The models were trained using the Adam optimizer with a learning rate of 3e-4, 0.1 gradient clipping on the Euclidean norm, and a weight decay of 0.05. They were trained with batch size 80, for a maximum of 10K steps. We used a polynomial learning rate scheduler with 1000 warm-up steps and a final learning rate of 3e-05. All the experiments were run on a single NVIDIA Titan X Maxwell 12GB GPU, taking around 24h for the training of the base model, and around 16h for boosted model. The DPO finetuning was done on the same machine with the trl library [^4] https://huggingface.co/docs/trl/en/index, using learning rate 5e-5, batch size 2, $\beta$ 0.1 and running it for 5 epochs, taking around 20min to finetune. During inference, we run all our models with 10 beams.
@@ Appendix H Additional analysis of cIE methods
SynthIE (constrained) SynthIE (unconstrained)
Unexhaustive 0.33 ± 0.09 0.23 ± 0.11
Incorrect related 0.36 ± 0.11 0.46 ± 0.13
Misclassified entity 0.09 ± 0.07 0.03 ± 0.03
Unrelated 0.60 ± 0.11 0.26 ± 0.11
Entity-centered 0.16 ± 0.08 0.34 ± 0.11
Entity/relation not in the KB 0.00 ± 0.00 0.72 ± 0.11
Table 4: Error analysis for SynthIE model in constrained and unconstrained mode on Wikipedia real text.
@@
As mentioned in Sec. 2, we suspect that errors that stem from the constrained and unconstrained decoding are complementary (see Fig. 1). To support this claim, we perform an analysis of the errors that occur during constrained and unconstrained decoding. We use the same error classification and Wikipedia data as in Sec. 5.3. We also add another type of error that only happens during the unconstrained decoding – Entity/relation not in the KB – which covers the case for which generated entities or relations are not present in the knowledge base. The results are shown in Table 4.